HUBBLE SPACE TELESCOPE DAILY REPORT #2354 PERIOD COVERED: 0000Z (UTC) 04/19/99 - 0000Z (UTC) 04/20/99 Daily Status Report as of 110/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed STIS/CCD/MA1/CCD 7432 (Morphological and Spectral Evolution of Galaxies) The Space Telescope Imaging Spectrograph (CCD and MA1) was used to make observations to study the morphological and spectral evolution of the galaxy 3C433-1. A library of ultraviolet images of normal galaxies at low redshifts is essential to our understanding of the properties and evolution of the high redshift universe. The observations were completed as planned, and no problems were reported. 1.3 Completed WF/PC-2 8053 (WF/PC-2 Cycle-7 Earth Flats -- Part II) The WF/PC-2 was used to take Earth flat calibration images. This proposal obtains sequences of Earth streak flats to construct high quality flat fields for the WF/PC-2 filter set. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Two Sets of STIS/CCD 7949 (CCD Dark and Bias Monitor -- Cont. to June 99) The Space Telescope Imaging Spectrograph (CCD) was used to conduct a CCD dark calibration and a bias monitoring program. It was also scheduled to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed STIS/CCD/MA2/WF/PC-2 7515 (UV-Optical Spectra and Imagery of the Bubble Galaxy NGC 7635) The Space Telescope Imaging Spectrograph (CCD/MA2) and the WF/PC-2 were used to acquire UV-optical STIS spectra and WFPC2 imagery of the wind-blown Bubble Nebula NGC 7635. The nebula's nearly spherical shell is the result of a recent { < 10^6 years} stellar mass-loss event and is the best young, clearly observed bubble available for study. The ionization front at the rim of the bubble will be examined, and the extent to which it is shock-driven and the scale of the photoevaporative flow off the face of the molecular cloud will be determined. A wave length and a dark calibration observation were also taken. The observations were completed as scheduled, with no reported problems. 1.6 Completed Two Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.7 Completed Two Sets of STIS/CCD 8091 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude, non-scripted, parallel observations. The observations were completed as planned, and no anomalies were noted. 1.8 Completed STIS/CCD 7361 (The Nature of Nuclear Activity in Nearby Galaxies) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the nuclear region of galaxy N4314. The high resolution spectra will be used to search for broad components in H-Alpha and for velocity gradients, indicative of gas rotation or other ordered motions. A wavelength calibration observation was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed WF/PC-2 7504 (Calibration of Nearby Type Ia Supernovae as Standard Candles: The Next Three-Year Step) The WF/PC-2 was used to make observations of parent galaxy NGC-4527 in order to determining its Cepheid distance. This is to continue the calibration of the Type 1a supernova as standard candles. The observations were completed as planned, and no porblems were reported. 1.10 Completed STIS/CCD/MA1/MA2 7301 (Studies of the Characteristics of the Interstellar Medium in the Carina Nebula) The Space Telescope Imaging Spectrograph (CCD, MA1 and MA2) was used to make observations of the star CPD-59D260 which is located in the Carina Nebula. It is proposed to carry out high-resolution ultraviolet spectroscopic observations of interstellar lines in the spectra of OB stars in the Carina Nebula, which will provide unique data on the interstellar medium associated with star cluster formation. A wavelength calibration image was also taken. The observations were completed as planned, and no anomalies were noted. 1.11 Completed WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed WF/PC-2 7619 (WF/PC-2 Cycle 7 Decontamination) The WF/PC-2 was used to conduct a Cycle 7 decontamination procedure. Ultraviolet-blocking contaminants were removed, and hot pixels cured, by warming the CCDs to +20C for 6 hours. A variety of internal images were taken before and after the decontamination in order to verify its success and to monitor the instrument health. The procedures were conducted as scheduled, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 5 Successful : 5 2.2 FHST Updates: Scheduled : 14 Successful : 14 2.3 Operations Notes: The SSR EDAC errors were cleared once using ROP SR-1. A NSSC-1 status buffer dump was performed per ROP NS-3. The Gyro-3 drift rate bias slope for the last seven-day interval is -7.95 arcseconds/hour/day (was -6.74). The slope for the last two-day interval is -4.39 arcseconds/hour/day (was -12.75). These slopes meet the limit of +/-10.8 arcseconds/hour/day for reinstallation of Gyro-3 into the control loop. PCS will continue to monitor the bias trend to verify long-term stability. Using ROP DF-7, the ephemeris and mag field tables were loaded, dumped, and compared. Per operations requests, the NICMOS FPA temperature sensor for detector 1 was enabled at 109/1059Z and at 109/1533Z, and the NICMOS FPA temperature sensor for detector 2 was enabled at 109/1103Z. The CCS-A HSTPS science data handler was down from 109/1936Z until 110/0342Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH